### v1

# WOMD-Reasoning: A Large-Scale Language Dataset for Interaction and Driving Intentions Reasoning

Abstract
We propose Waymo Open Motion Dataset-Reasoning (WOMD-Reasoning), a language annotation dataset built on WOMD***Waymo Open Motion Dataset and WOMD are trademarks of Waymo LLC, and are used here by permission., with a focus on describing and reasoning interactions and intentions in driving scenarios. Previous language datasets primarily captured interactions caused by close distances. However, interactions induced by traffic rules and human intentions, which can occur over long distances, are yet sufficiently covered, despite being very common and more challenging for prediction or planning models to understand. Therefore, our WOMD-Reasoning focuses extensively on these interactions, providing a total of 409k Q&As for varying types of interactions. Additionally, WOMD-Reasoning presents by far the largest Q&A dataset on real-world driving scenarios, with around 3 million Q&As covering various topics of autonomous driving from map descriptions, motion status descriptions, to narratives and analyses of agents’ interactions, behaviors, and intentions. This extensive textual information enables fine-tuning driving-related Large Language Models (LLMs) for a wide range of applications like scene description, prediction, planning, etc. By incorporating interaction and intention language from WOMD-Reasoning, we see significant enhancements in the performance of the state-of-the-art trajectory prediction model, Multipath++, with improvements of 10.14% in $MR_{6}$ and 6.90% in $minFDE_{6}$, proving the effectiveness of WOMD-Reasoning. We hope WOMD-Reasoning would empower LLMs in driving to offer better interaction understanding and behavioral reasoning. The dataset is available on https://waymo.com/open/download.
Figure 1: A typical example of interaction descriptions and reasoning in WOMD-Reasoning. In this case, WOMD-Reasoning successfully captures the traffic rule-induced interaction between the ego agent and agent # 0, attributing it correctly to the stop signs.
1 Introduction
Large-guided driving datasets and models have drawn much attention in autonomous driving. For example, methods based on large language models (LLMs) [27, 1] have shown effectiveness in improving explainability, controllability, and performance of driving tasks [25, 5, 10, 14, 33]. Various language datasets [25, 19, 12, 23, 14, 11] have been proposed for fine-tuning these LLMs. However, due to the complexity and variability of driving scenarios and driver behaviors, existing datasets often focus on specific scenes without covering critical cases, e.g., reasoning about interactions and human intentions, which is vital for the advancement of complex decision-making processes in autonomous driving systems [29, 36, 37]. For example, in BDD-X [12], interactions are often limited in spatial relationships, e.g., ‘The car moves back into the left lane because the school bus in front of it is stopping.’, while in an example, DriveLM [25] attributes interaction to ‘keep a safe distance’ without further detailed explanations. As a result, most existing works have drawbacks, either insufficient in quantity or only covering interactions induced by proximity, as shown in our analysis in Table 1.
Considering the critical importance and difficulty of interaction modeling for autonomous driving [8, 21], it is essential to incorporate comprehensive interactive and behavioral reasoning for language-based driving. Real-world interactions are often driven by human knowledge and behavior, such as traffic rules, signals, and intentions like overtaking. However, these interactions are rarely covered in previous datasets, leading to suboptimal performance of LLMs fine-tuned on these datasets. Additionally, most existing language datasets are not large enough to support LLM fine-tuning across diverse tasks such as scene description, prediction, and planning.
Therefore, we build WOMD-Reasoning, a large-scale language Q&A dataset based on WOMD [9], with a focus on interaction descriptions and reasoning. To incorporate interactions induced by traffic rules and human intentions, we build an automated data curation pipeline based on prompting ChatGPT [1, 18] with a set of well-designed prompts. Together with a rule-based translator to convert motion data into language, we create an automated pipeline for generating the language dataset, achieving an overall accuracy rate of over 90%. We then manually verify and correct a subset to provide a high-quality set suitable for supervised fine-tuning (SFT) with LLMs. To the best of our knowledge, WOMD-Reasoning is the currently largest real-world language dataset for driving.
The resultant WOMD-Reasoning dataset successfully covers interactions induced by traffic rules and human intentions with 409k Q&As, like those shown in Figure 1. As indicated in Table 1, it contains the largest number of total Q&As and interaction-specific Q&As, supporting a wide range of applications including scene description, prediction, and planning. We empirically find that language in WOMD-Reasoning can significantly enhance the performance of vehicle trajectory prediction tasks. Qualitative results indicate a good consistency between these performance improvements and the corresponding language inputs. Our contributions can be summarized in the following:
  • We propose WOMD-Reasoning, a language dataset centered on interaction descriptions and reasoning. It provides extensive insights into critical but previously overlooked interactions induced by traffic rules and human intentions.
  • We develop an automatic language labeling pipeline, leveraging a rule-based translator to interpret motion data into language descriptions, and a set of manual prompts for ChatGPT to generate Q&A pairs. As a result, WOMD-Reasoning emerges as, to the best of our knowledge, the largest language dataset for real-world driving, comprising approximately 3 million Q&A pairs supporting scene description, prediction, and planning tasks.
  • We validate the effectiveness of WOMD-Reasoning in vehicle trajectory prediction tasks. By training Multipath++ on WOMD+WOMD-Reasoning, significant performance improvements have been observed, e.g., 10.14% in $MR_{6}$ and 6.90% in $minADE_{6}$. Our findings suggest a notable correlation between the language input and the improvements in prediction accuracy, proving the effectiveness of interaction narratives in WOMD-Reasoning.
Table 1: Comparison between WOMD-Reasoning and Previous Real-world Language Datasets for Driving Scenarios. Compared to previous datasets, WOMD-Reasoning covers comprehensive kinds of interactions, with a significantly higher size supporting a wide range of potential applications
Dataset Data Source Statistics Interactions Applications
Total Scenes Total Q&As Interaction Q&As Distance-induced Traffic Rule-induced Human Intention-induced Scene Descriptions Motion Prediction Motion Planning
nuScenes-QA[19] nuScenes[2] 34k 460k 0
BDD-X[12] BDD[32] 26k 26k 26k
DriveLM[25] nuScenes[2] 5k 443k $\approx$199k
DRAMA[14] DRAMA[14] 18k 102k <18k
WOMD-Reasoning WOMD [9] 63k 2,940k 409k
2 Related Works
Application of Language Modal in Autonomous Driving. Language is widely utilized in autonomous driving models, as incorporating language into implicit features could enhance the explain-ability [34, 22, 26, 3, 15]. For example, Kuo et al. [13] employ language features in the trajectory prediction model, ensuring both explainability and consistency between language and predictions. Similarly, Zhong et al. [38] use language input and large language models (LLMs) to generate conditions in diffusion models, thereby controlling the generation of driving scenarios. Recently, general driving agents have been developed [33, 5, 25, 16], aiming to integrate all functions—perception, prediction, and planning—into a single language agent. However, we observe that their performance in interaction analysis is not entirely satisfying. This shortcoming is largely attributed to the lack of interaction analysis in language datasets specific to driving scenarios.
Language Datasets in Autonomous Driving. Language datasets for driving scenarios have recently been developed to support related LLM-based work. Due to the controllability of simulations, several simulation-based datasets have been created [5, 25]. However, simulations often fail to capture real-world interactions. To address this, recent studies [12, 23, 14] incorporate human labeling of real-world driving scenarios to ensure the inclusion of genuine interactions. Nevertheless, obtaining high-quality human labels is labor-intensive and costly, limiting the size and coverage of these datasets. Moreover, interactions driven by traffic rules and human intentions have often been overlooked. To streamline the labeling process, some studies [7, 17, 19, 25] have generated rule-based labels, although these typically cover only basic language elements like scene descriptions, leaving interaction analysis to human labelers. To minimize human labor, we first employ manual rules and then utilize ChatGPT-4 [18] to build our dataset automatically. The resulting dataset, WOMD-Reasoning, is rich in interaction details, particularly those stemming from traffic rules and human intentions. The designed automated data-curation pipeline enables the dataset to be significantly larger. We also create a high-quality human-verified subset that can be used for supervised fine-tuning. Preliminary human evaluation provides positive feedback with an overall accuracy of over 90%.
3 Method
Building language datasets would involve intense human labor, which is one of the main reasons why previous datasets have insufficient interaction analysis and a very limited overall size [24]. To reduce human labor while obtaining reasonably useful data on vehicle interactions, we propose a fully automatic data-curation pipeline to label the motion dataset with language: We first develop a rule-based program to interpret the motion dataset, which contains trajectories and the HD map, into language; and then build a set of prompts to utilize the reasoning abilities of ChatGPT-4 [18] to generate interaction analysis and reasoning and organize the results into the target Q&A format.
3.1 Automatic Translation of Driving Scenarios into Language Descriptions
ChatGPT cannot effectively process raw data from motion datasets, as it is mainly trained on textual information rather than motion data. Therefore, we first translate the motion information into textual descriptions by a rule-based program. Accurate motion translation is critical to the quality of the dataset, as it serves as the only input observable to GPT. Ensuring thorough and well-structured translations is essential for GPT to perform accurate interaction analysis and reasoning.
To ensure enough information for analyzing interactions involving the ego agent, we translate the motion data into an ego-centric description of the traffic scenario. Specifically, before describing the ego agent status, we first narrate the map environment around the ego agent, including information about related intersections, lanes, stop signs, crosswalks, speed bumps, etc. The above information provides a comprehensive spatial guidance that facilitates describing the motion of either ego or other agents accurately. Then, we describe the status of the ego agent in the scenario, including its motion status (e.g., the velocity, acceleration, etc.), and the positional status (e.g., the lane it occupies and its position related to the intersection center). Furthermore, its related traffic light and signal information is also summarized in the description. Besides the ego agent, the depiction of other agents, including other vehicles, bicycles, and pedestrians is also crucial for GPT to analyze the interactions. Therefore, we then describe every other agent in the scenario in the same way as we describe the ego agent. We additionally include the positional relation between the other agent and the ego one to support the identification of every agent’s position.
Utilizing the methods above, we are able to translate motion datasets into languages. Due to the limitations of computational resources of GPT, we only choose to translate those scenarios that have ”objects_of_interest” labels in WOMD, which indicates confirmed significant interactions happening in the scenario. We build two subsets of WOMD-Reasoning with the same setting: the training set is built on the training set of WOMD, while the validation set is built on the interactive validation set of WOMD. In total, we translate 63k scenarios into language. An example of the translation can be found in Table 5 in the Appendix.
3.2 GPT-based Interaction Analysis
With the language translation of WOMD as input, we further design a set of prompts to utilize ChatGPT-4 [18] to build the Q&A dataset, taking advantage of its abilities in analyzing interactions. The prompts come with 4 main parts. (1) System Prompt, which lets the GPT know its role and the format of its input, i.e. the language translation. (2) Responsibility Prompt, which describes the responsibilities of the GPT, including the questions it should ask and answer, as well as the output format it should use. This will be talked about in detail in the next paragraph. (3) Rules Prompt, which contains global rules to guide the output and the analysis. (4) In-context Prompt, which provides a few human-written examples of input-output pairs to perform in-context learning. An example of each part of the prompts can be found in Table 6, 7, 8, and 9 respectively in the Appendix.
The questions in the responsibility part of the prompt decide what would be present in the generated language Q&A dataset. To provide a comprehensive language dataset, and to present interaction analysis as accurately as possible, we organize the questions into a chain to enable a well-designed chain-of-thought [30, 31, 6]: First, we ask questions about the map environment, the ego and other agents’ motion. These questions would let the GPT provide Q&As useful for fine-tuning scene description agents while enhancing the GPT’s understanding of the input information at the same time. Then, we ask questions about the interactions that occur in the observable future period. To include traffic rule-induced and intention-induced interactions, we ask the GPT to think about the interaction between each pair of agents by answering a sequence of questions: Q1: Are the two agents vehicles, and are they close to each other with traffic rules or signals governing their movements? If the answer is yes, we request GPT to give an analysis of their yielding relations according to the traffic rules. Otherwise, we ask Q2: Are the two agents vehicle and pedestrian respectively with intention conflicts? If so, the vehicle generally should yield to the pedestrian for safety. If both questions do not help find the interaction relations, we come to Q3: Does the scene show patterns of certain intention-induced interaction? We provide descriptions of a few common interactions like overtaking, following, etc. GPT can choose the best fit from these patterns to try to cover more interaction types. Based on this pipeline, we are able to provide interaction analysis with rich information on traffic rule-induced and human intention-induced interactions. And after that, finally, we are able to wrap up the interactions between the ego and all other agents, to provide the intention of the ego agent.
Our analysis process is performed on Microsoft Azure ChatGPT-4-Turbo API, which costs around 12,750 USD. The details of WOMD-Reasoning will be provided in the following paragraph.
Figure 2: An Additional Example of Traffic Rule-based Interaction and Intention in WOMD-Reasoning. In this scenario, WOMD-Reasoning completely covers the traffic light-controlled yielding interaction between the ego agent and agent # 1.
Table 2: Summary of dataset split and the quantitative distribution across different categories in the WOMD-Reasoning dataset.
Categories Map Ego Agent Other Agents Interactions Intentions Total Scenes
Training Set 188k 268k 1,635k 287k 52k 2,430k 52k
Validation Set 41k 58k 341k 59k 11k 510k 11k
Total 229k 326k 1,976k 346k 63k 2,940k 63k
Per Scene Ave. 3.58 5.09 30.88 5.41 0.98 45.95 -
4 Dataset
In this section, we provide detailed information about WOMD-Reasoning. We first present the organization and statistics of the dataset, then provide quality analysis based on case studies and preliminary human evaluations. Finally, we compare our dataset to previous ones to show the strengths of WOMD-Reasoning in diverse interactions, huge size and wide potential applications.
4.1 Dataset Details
Table 2 provides the statistics of Q&As in WOMD-Reasoning dataset, as well as in every Q&A class. Our Q&As can be classified into five categories: Map Environment, which includes the existence and category of intersections, existence and places of stop signs and crosswalks, and count of lanes, etc. Ego Agent’s Motion Status and Other (Surrounding) Agents’ Motion status, which describe each agent’s speed, acceleration, direction, related traffic light and sign, and relative position to the intersection center as well as to the ego agent. Interaction, which includes a summary of interactions between each surrounding agent and the ego agent, and the reasoning behind such interactions. Intention, which sums up all interaction information to give the predicted intention of the ego agent, considering its responses to all the interactions. The description of the map environment and agents’ motion status in the dataset are good for training language scene describers, while the interaction and intention parts can be used for training language-based prediction and planning agents.
We claim that to the best of our knowledge, WOMD-Reasoning is the largest real-world language dataset for driving. As shown in Table 1, we provide roughly 3 million Q&A pairs for 63k scenes in WOMD, which is 6-113 times the total Q&As of the datasets we compare to. Furthermore, WOMD-Reasoning contains 346k unique traffic rule-induced and intention-induced interaction Q&As, as well as 63k comprehensive intention predictions of the ego agent containing responses to the interactions, which rarely exist in previous datasets. We will show these features through examples and statistics in the following section.
4.2 Dataset Quality
To testify to the quality of WOMD-Reasoning, we show examples from it containing rich traffic rule-induced and human intention-induced interactions.
WOMD-Reasoning contains traffic rule-induced interaction information Firstly, in Figure 1, we show a traffic scenario of an intersection controlled by the stop sign. In this case, the agent #0 is quite far away from the ego agent at the starting (current) moment, thus their interactions would rarely be considered in previous datasets, as they mainly cover distance-induced interactions. However, the two agents do have a significant interaction: Controlled by the stop signs, the agent #0 has to yield to the ego agent, even though the ego agent is still far from the intersection. This case vividly proves the remarkable importance of the traffic rule-induced interactions we cover.
Another example is a traffic light-controlled intersection shown in Figure 2. Similar to the previous one, agent #0 is initially not close to the ego agent. However, thanks to our analysis of traffic rules, our language dataset effectively captures the interaction that the agent # 1 would yield to the ego agent who is turning left and having right-of-way due to an arrow green light. This further proves that traffic rules are essential in determining the interactions.
Figure 3: A demonstration of Q&As in each part of WOMD-Reasoning and an example for human intention-induced interaction. We show Q&As in all 5 categories regarding the scenario while demonstrating language analysis of overtaking, a human intention-induced interaction in WOMD-Reasoning.
Table 3: Preliminary Human Evaluation of the Dataset. Note that we ask evaluators to comment on whether main part of the answer is correct, where minor errors would not be considered as wrong.
Q&A Class Map Ego Agent Other Agents Interactions Intentions
Total Q&As 38 57 339 57 10
Average ”Accuracy” Rate 80.70% 95.32% 93.51% 94.73% 90.00%
Standard Error 6.08% 2.03% 3.08% 3.51% 10.00%
WOMD-Reasoning contains human intention-induced interaction information Beside traffic rules, our dataset also covers human intention-induced interactions. We provide an additional example to show this, as well as to demonstrate a whole picture of each section of the dataset. In Figure 3, WOMD-Reasoning provides fruitful information on the human intention-induced interaction - overtaking, based on pattern recognition abilities built in the GPT. We also show 1-2 Q&As for each category of the dataset for completeness.
To grab a sense of the accuracy of WOMD-Reasoning, we also execute a preliminary human evaluation. Consent has been obtained from all people interviewed. Note that here we ask evaluators to think a Q&A pair as correct if the correct main answer to the question is included in the answer, so the main answer is reasonably based on the human evaluators’ interpretation of the traffic data. This also does not exclude minor errors (like errors in detailed numbers or very far-away agents), since such errors are hardly distinguishable by the human evaluators. We invite 3 people to judge all Q&As from 11 consecutive scenes of the dataset, which contains 501 Q&As in total. The results are shown in Table 3. We conclude that considering our dataset is fully automatically generated, it achieves a satisfying accuracy, or at least it would significantly reduce the human labor necessary to build a more reliable dataset later. Further human verification of the dataset is still ongoing.
4.3 Comparison to Existing Datasets
Comparing WOMD-Reasoning with existing datasets, as shown in Table 1, we are the largest real-world dataset, in terms of the total scenes, total Q&As, and in the total interactions included. Also, WOMD-Reasoning supports Q&As in scene descriptions, prediction and planning, making it suitable for training LLMs to perform nearly all major tasks in autonomous driving. Diving even deeper into the interaction reasoning in these datasets, we summarize that most interactions in previous datasets are caused by a very near distance. However, as we show in Sec. 4.2, many interactions happen when the two counterparts are initially far away, but they establish interactions by following the traffic rules or by human intentions. Our WOMD-Reasoning contains a wealth of this interaction information, as shown in Figure 4, the vocabulary statistics of the interaction and intention part of our dataset.
Figure 4: Selected vocabulary statistics of WOMD-Reasoning. We stat vocabularies strongly related to traffic rule-induced and human intention-induced interactions in the interaction and intention part of WOMD-Reasoning, illustrating that our dataset contains an abundance of such interaction descriptions and reasoning.
5 Dataset Effectiveness Validation with Trajectory Prediction Task
In this section, we examine the effectiveness of our language-based interaction analysis with vehicle trajectory prediction tasks. We build a very simple network combining T5 text encoder [20] and Multipath++ [28], seeing significant performance boost when the language is introduced.
Figure 5: Effect of language dataset on trajectory prediction. (a) shows the generated language leads the ego agent to be cautious on the incoming car when turning left, while (b) shows the generated language leads the ego agent not to be too cautious when the other agent is yielding to it.
5.1 Method
Our pipeline is simple, designed for demonstrating the effectiveness of the language dataset. We first follow  [35, 4] to use T5 text encoder [20] to encode the language. Specifically, to show the influence of interaction and intention data, we encode these two parts of language data, followed by a few MLPs to reduce their dimensions. Then, we allow the features of the ego and other agents in Multipath++ pipeline to cross-attend to representations of the ego’s intention language and the other agents’ interactions language respectively. The model is trained end-to-end, with the T5 encoder frozen. The pipeline is illustrated in Fig 6 in Appendix, and implementation details are listed in Sec A.1
Due to the restriction of computational resources, we have not fine-tuned LLMs to generate the interaction language analysis. Therefore, our experiment is only an oracle model method, where the interaction language is directly taken from the dataset with information leaks, i.e. the interaction analysis are imported from the dataset directly. However, this can still prove whether the languages in our dataset are useful, though it remains a question whether it can be generated successfully after fine-tuning the LLM on our dataset.
We perform experiments on both the whole WOMD and part of WOMD with language from WOMD-Reasoning. Note that to ensure consistency, we test the model trained on the whole WOMD training set on the whole WOMD interactive evaluation set, while testing the model trained on the language available part of training set on language available part of evaluation set.
5.2 Quantitative Results
As shown in Table 4, with the introduction of the language, the performance of the trajectory prediction model is substantially enhanced. When training and evaluating on the part of WOMD having language, the language modal helps to reduce the Miss Rate (MR) by $10.14\%$ and the minFDE by $6.90\%$. These significant improvements strongly support the effectiveness of the interaction analysis language in downstream tasks like trajectory predictions.
Table 4: Effectiveness of WOMD-Reasoning on the Prediction Task.
Motion Dataset Metric NOT Using Language Using Language
Whole WOMD $MR_{6}(\downarrow)$ 10.91 10.59
WOMD $\cap$ WOMD-Reasoning $MR_{6}(\downarrow)$ 11.44 10.28
WOMD $\cap$ WOMD-Reasoning $minFDE_{6}(\downarrow)$ 1.16 1.08
5.3 Qualitative Results and Analysis
To thoroughly understand the role that language dataset plays in the language-guided trajectory prediction model, and to verify the correspondence between the language data and the prediction result, we illustrate a few scenarios whose prediction is improved with language. In Figure 5 (a), with the help of language noticing that the ego agent should wait for the traffic light before turning left, the ego agent would be more cautious and its predicted conflict with agent #0 is significantly reduced. Reversely in Figure 5 (b), under the language input saying the ego agent does not need to yield to any other agent, the predicted trajectories do not slow down, thus being closer to the ground truth. Both cases suggest that the prediction results improve consistently with the language.
6 Conclusion
We build WOMD-Reasoning, a language annotation dataset upon WOMD to support LLMs for analyzing and reasoning interactions. Specifically, we focus on interactions induced by traffic rules and human intentions, such as interactions governed by traffic lights and signs, and interactions like overtaking and following, which are both rarely covered in previous datasets. Besides, our WOMD-Reasoning provides 3 million Q&A pairs, which makes it the largest real-world language dataset in autonomous driving to the best of our knowledge, covering a range of topics from scene descriptions, prediction, to planning. By utilizing our proposed dataset, we significantly improve the performances of the trajectory prediction model Multipath++, where we see matches between the language and the prediction improvements, proving the effectiveness of our interaction analysis for downstream tasks.
Limitations. Although WOMD-Reasoning demonstrates its effectiveness in case studies, human evaluations, and downstream prediction tasks, we only manually verified a small subset. With human verification ongoing, the dataset will achieve even greater reliability. Also, WOMD-Reasoning is subject to the limitations of WOMD. For instance, the raw camera data is not yet publicly available, which restricts the use of WOMD-Reasoning in video-qa. As WOMD continues to release additional modalities such as camera features and lidar data recently, these limitations are expected to be naturally addressed.
Broader Impact. Our work is of broad interest to the natural language processing (NLP) and autonomous driving communities. The proposed approach has no new ethical or social issues on its own, except those inherited from NLP or autonomous driving.
7 Acknowledgements
This work was supported in part by Berkeley DeepDrive, NSF Grant. 2235013, and the Microsoft Accelerate Foundation Models Research Initiative.
Appendix A Appendix
In this supplementary material, we first show the pipeline we use to introduce language into Multipath++ trajectory prediction model, as well as the training details of Multipath++ on WOMD-Reasoning. Then, we show an example of the interpreted language of the motion data. We also revisit the prompts to build the language dataset. We will make the dataset and codes public.
A.1 Pipeline for Introducing Language into Multipath++
Figure  6 presents the pipeline of introducing language into the Multipath++ trajectory prediction pipeline. We use a batch size of 128 and a learning rate of 1e-4 for Multipath++. For training or evaluating on the whole dataset, some scenarios do not have language, so we input "No captions available for this scene." as the language input. Also even for scenarios with language, language is not available for every agent, so we use "No captions available for this agent." for those agents without language inputs. The t5-v1_1-xxl model is used as the language encoder.
Figure 6: Pipeline of introducing language data into Multipath++
A.2 Example of Translated Motion Scenario
Table  5 presents a part of the interpreted motion scenario by the rule-based program we design.
Descriptions on map environment:
The ego agent is heading towards intersection. The intersection center is 18.0 meters in front of the ego agent, and is 17.0 meters on the left of the ego agent. The intersection is a 4 way intersection. Descriptions on ego agent:
The ego agent is on the 1 lane from the right, out of 3 lanes. Its current speed is 8 m/s. It is accelerating. Traffic Light for the ego agent is red. The ego agent is approaching a crosswalk 3 meters ahead. Descriptions on surrounding (other) agents:
Surrounding agent # 0 is a vehicle. It is 23 meters in front of the ego agent, and is 17 meters on the left of the ego agent. It is heading right of the ego agent. Its current speed is 6 m/s. It is accelerating. It is in the intersection. It is 29 meters away from the ego agent. It is approaching a crosswalk 3 meters ahead. Surrounding agent # 2 is a vehicle. It is 11 meters on the right of the ego agent, and is 0 meters behind the ego agent. It is heading left of the ego agent. It is not moving. It is on the same side of the intersection as the ego agent. It is 34 meters away from the intersection center.
$\cdots$
Descriptions on agents after 3 seconds:
The following is the description of the ego and surrounding agents after 3.0 seconds: Surrounding agent # 0 will be 8 meters in front of the ego agent, and will be 5 meters on the left of the ego agent. It will be heading right of the ego agent. Its speed will be 7 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 18 meters away from the intersection center. The ego agent will be 13 meters in front of the current place, and will be 2 meters on the right of the current place. It will be heading in the same direction as the current moment. Its speed will be 3 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 20 meters away from the intersection center. Surrounding agent # 2 will be 13 meters on the right of the ego agent, and will be 9 meters behind the ego agent. It will be heading left of the ego agent. It will not be moving. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 34 meters away from the intersection center.
$\cdots$
Table 5: An example of the language translation of motion data.
A.3 Prompts for GPT
A.3.1 System Prompt
Table  6 presents the system prompt part of our prompt set.
System prompt for GPT:
You are an AI assistant to analyze real-world driving scenes. You are provided with the detailed information (formated as: [start of the input] the detailed information [end of the input]) of the real-world driving scenario, which primarily includes information about the ego vehicle (denoted as ’Ego Agent’), other vehicles, cyclists or pedestrians (denoted as ’Surrounding agent #n’), intersections and road lanes. Note that the input includes the current situation as well as the situation in a future moment for your reference. Unless you are told to guess, only provide information you are certain according to the input. The input is complete. Any agents or other things not mentioned in the input does NOT exist, and you should NOT guess what would happen if they exist.
Table 6: System prompt for GPT.
A.3.2 Responsibility Prompt
Table  7 presents the responsibility prompt part of our prompt set.
Responsibility Prompt for GPT:
You are responsible for the following responsibilities: <First>, print "[Env QA]". Generate Q&A on all the scenario description inputs about the scenario environment. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" The Q&A should cover ALL information about the CURRENT moment in the input description. No information in the provided future moment should be used. Your questions should include all the questions in the following categories except the answer is not mentioned in the input: (1) About the intersection: its existence? Its type? (2) About the lanes: How many lanes on ego’s side? (3) About stop signs: its existence? where is it? which direction is it for?
$\cdots$
<Second>, print "[Ego QA]". Generate Q&A on all the scenario description inputs about the the ego agent. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" The Q&A should cover ALL information about the CURRENT moment in the input description. No information in the provided future moment should be used. Your questions should include all the questions in the following categories: (1) About its motion: its speed? its motion status? (2) About its position: its position? its lane? (3) About its direction: its facing direction? its turning direction?
$\cdots$
<Third>, print "[Sur QA]" and generate Q&A on all the scenario description inputs about surrounding agents. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" The Q&A should cover ALL information about the CURRENT moment in the input description. No information in the provided future moment should be used. For each surrounding agent, your questions should include all the questions in the following categories except the answer is not mentioned in the input: (1) About itself: its type? (2) About its motion: its speed? its motion status? (3) About its position: its position to the ego agent? its position to the intersection? its lane?
$\cdots$
<Fourth>, start a new paragraph and print "[Int QA]". For each surrounding agent, start a new paragraph and provide one Q&A for the interactions between it and the ego agent. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" Here [Question content] should have the meaning "What interactions will happen between surrounding agent #n and the ego agent?" without any additional information like "considering what" or "given what", but you must use diverse ways to put it. (i.e. ask in different ways for different surrounding agents) When providing [Answer content], you should use the descriptions for both the current moment and the future moment, and analyze based on them. Pay attention to the special situations where you should consider that there is an interaction:
$\cdots$
<Fifth>, for the ego agent, start a new paragraph, print "[Intention]", and provide one Q&A for the intention of the ego agent. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" Here [Question content] should have the meaning "What will be the intention of the ego agent?", but you must use diverse ways to put it. When providing [Answer content], and think about following questions in the time period from now to the provided future moment based on given description: (1) What is the ego agent’s intention if no other agents exist? (2) If applicable, what actions does the ego agent take to respond to traffic lights, stop signs, crosswalks, speed bumps and traffic rules? If some traffic controls are not applicable, do not output on that. (3) In each interaction involving the ego agent, what actions do the ego and surrounding agent take to respond to each other? If one agent does not have interaction with the ego agent, exclude it in your answer.
$\cdots$
Remember to use diverse ways to put the [Answer content]. Do not include any numbers regarding speeds or positions in your [Answer content].
Table 7: Responsibility prompt for GPT.
A.3.3 Rule Prompt
Table  8 presents the rule prompt part of our prompt set.
Global rules prompt for GPT:
Here are some global rules that your need to follow: 1. Please provide the summary and analysis of the driving scenario according to your responsibilities. 2. Never use bullet points. Be concise and compact. 3. Give natural language, no parenthesis. 4. Give the results in the order of the responsibilities. Always start a new paragraph for each responsibility, but one responsibility can take several paragraphs. Do not combine multiple responsibilities together. Finish one responsibility before starting another. 5. Do not show the intersection number, lane number or the ego agent’s number. The surrounding agents’ numbers need to be shown. 6. Output in full. Never omit anything regardless of any reason. You are allowed to give very long outputs. 7. Remember that all the agents and traffic controls in the scene have been provided, so if any agents or traffic controls are not mentioned in the scene, they do not exist. You don’t need to consider them or give any output about them. Therefore, do not output words like "not mentioned" because this is inaccurate. 8. The traffic light info is for the current moment only. The future traffic light info is unknown. 9. Any action caused by traffic controls like lights, stop signs or crosswalks may also constitute yielding interactions. Traffic controls are only an indication of right-of-way, but any actions that are caused by the traffic controls are still considered interactions. For example, if a car is waiting for the red light or a stop sign while someone other cross the intersection with green light, the car is still considered yielding to the other car. 10. When analyzing the future, talk about the period between the current and future moment provided, not the two moments.
Table 8: Rule prompt for GPT.
A.3.4 In-context Prompt
Table  9 presents the in-context prompt part of our prompt set.
Context prompt for GPT:
Here is one given example that you can refer to to help you better understand you responsibility <Fourth> and <Fifth>: For the input case of: [start of input] The ego agent is in intersection. The intersection center is 10.0 meters in front of the ego agent, and is 1.0 meters on the left of the ego agent. The intersection is a 4-way intersection. It is going straight. It is exiting the intersection. It ’s current speed is 6 m/s. It is accelerating. The ego agent is approaching a stop sign 1 meters ahead. There are 4 stop signs in the intersection. The ego agent is approaching a crosswalk 1 meters ahead. Surrounding agent # 0 is a vehicle. It is 4 meters on the left of the ego agent, and is 2 meters behind the ego agent. It is heading the opposite direction as the ego agent. Its current speed is 9 m/s. It is accelerating. It is departing from the intersection. It is on the same side of the intersection as the ego agent. It is 12 meters away from the intersection center.
$\cdots$ [end of input]
you may give the following analysis: [Int QA] [Q] What kind of interaction can be expected between the ego agent and surrounding agent #0 given their positions and directions? [A] Surrounding agent #0 will have no interaction with the ego agent as their intended paths have no conflicts.
$\cdots$ [End Int QA]
[Intention] [Q] What will the ego agent aim to do in the upcoming moments? [A] The ego agent intends to continue exiting the intersection. There is a stop sign on its side but it has already stopped and started afterwards. Surrounding agent #1 will yield to it so the ego agent does not need to respond. Surrounding agent #3 will follow it but the ego agent does not need to respond as well. Therefore, the ego agent will continue in the intersection and finish its left turn. [End Intention]
Here is another given example that you can refer to to help you better understand you responsibility <Fourth> and <Fifth>: [start of input] The ego agent is heading towards intersection. The intersection center is 28.0 meters in front of the ego agent, and is 3.0 meters on the left of the ego agent. The intersection is a 3 way intersection. The ego agent is on the 2 lane from the left, out of 3 lanes. It ’s current speed is 0 m/s. It is decelerating. Traffic Light for the ego agent is red 9.6 meters ahead. The ego agent is approaching a crosswalk 3 meters ahead. Surrounding agent # 0 is a vehicle. It is 29 meters in front of the ego agent, and is 4 meters on the right of the ego agent. It is heading left of the ego agent. Its current speed is 6 m/s. It is accelerating. It is in the intersection. It is 29 meters away from the ego agent. The ego agent is at a speed bump.
$\cdots$ [end of input]
you may give the following analysis: [Int QA] [Q] What kind of interaction can be expected between the ego agent and surrounding agent #0? [A] The ego agent will yield to surrounding agent #0 because they are both approaching the intersection and the traffic light is red for the ego agent, indicating that the ego agent must stop. [End Int QA]
[Intention] [Q] What will the ego agent aim to do in the upcoming moments? [A] The ego agent intends to enter the intersection, but it has to stop at the red traffic light and wait for it to change to green before proceeding through the intersection. It will also yield to surrounding agents #0 who is already in the intersection. Therefore, the ego agent will wait until the lights turn green and the surrounding agent #0 passes before entering the intersection. [End Intention]
$\cdots$
Table 9: In-context prompt for GPT.