### v2

# WOMD-Reasoning: A Large-Scale Dataset and Benchmark for Interaction and Intention Reasoning in Driving

Abstract
We propose Waymo Open Motion Dataset-Reasoning (WOMD-Reasoning), a comprehensive large-scale dataset with 3 million Q&As built on WOMD focusing on describing and reasoning interactions and intentions in driving scenarios. Existing language datasets for driving primarily capture interactions caused by close distances. However, interactions induced by traffic rules and human intentions, which can occur over long distances, are yet sufficiently covered. To address this, WOMD-Reasoning presents by far the largest multi-modal Q&A dataset on real-world driving scenarios, covering a wide range of driving topics from map descriptions and motion status descriptions to narratives and analyses of agents’ interactions, behaviors, and intentions. We further introduce Motion-LLaVA, a motion-language model fine-tuned on the proposed dataset with robust interaction reasoning capabilities. We benchmark its performance across various configurations including different input modalities, reasoning techniques, and network architectures. The robust, diverse, and multi-modal nature of WOMD-Reasoning highlights its potential to advance future autonomous driving research and enable a broad range of applications. The dataset and its vision modal extension are available at https://waymo.com/open/download, and the codes & prompts to build it are available at https://github.com/yhli123/WOMD-Reasoning.
Figure 1: Two examples of interaction descriptions and reasoning in WOMD-Reasoning dataset. (a) This example captures the traffic rule-induced interaction between the ego agent and agent #0, attributing it correctly to the stop signs. (b) This example shows the traffic light-controlled yielding interaction between the ego agent and agent #1. The front-view visualization is created by importing the traffic scenario into MetaDrive simulator [24].
1 Introduction
Language-guided driving datasets and models have drawn much attention in autonomous driving. For example, methods based on large language models (LLMs) [50, 4] improve explainability, controllability, and performance of driving tasks [7, 14, 30, 57, 23, 2, 36, 26, 9, 29, 60, 16, 11, 17, 53].
Various language-based datasets [45, 37, 20, 41, 30, 18, 65, 48, 2, 36, 26, 29] have been proposed to fine-tune these LLMs. However, due to the complexity and variability of driving scenarios and driver behaviors, existing datasets often focus on describing specific scenes without reasoning about the interactions, like traffic rules and human intentions behind them, which are vital for the safety-critical decision-making in autonomous driving systems [54, 62, 63]. For example, in BDD-X [20], interactions are often limited in spatial relationships, e.g., ‘The car moves back into the left lane because the school bus in front of it is stopping.’; while in another example, DriveLM [45] attributes interaction to ‘keep a safe distance’ without further detailed explanations. As a result, most existing works are either insufficient in quantity or only cover interactions induced by proximity, as shown in our analysis in Table 1.
Considering the critical importance and difficulty of interaction modeling for autonomous driving [12, 39], it is essential to incorporate comprehensive interactive and behavioral reasoning for language-based driving. Real-world interactions are often driven by human knowledge and behavior, such as traffic rules, signals, and intentions like overtaking. However, these interactions are rarely included in previous datasets, leading to suboptimal performance of LLMs fine-tuned on these datasets. Additionally, most datasets in driving are not large enough to support multi-modal fine-tuning across diverse tasks such as scene description, prediction, and planning.
This work introduces WOMD-Reasoning dataset, a large-scale multi-modal dataset centered on language Q&As based on WOMD [13], with a specific focus on interaction descriptions and reasoning. To incorporate interactions induced by traffic rules and human intentions, we build an automated data curation pipeline by prompting ChatGPT [4, 34] with a set of well-designed prompts. Together with a rule-based translator to convert motion data into language, we create an automated pipeline for generating the language dataset and achieve an approximate accuracy rate of over 90%. We then manually verify and correct a subset to provide a high-quality set suitable for supervised fine-tuning (SFT) with LLMs. To the best of our knowledge, WOMD-Reasoning is currently the largest real-world language dataset for driving. WOMD-Reasoning dataset covers interactions induced by traffic rules and human intentions with 409k Q&As, like those shown in Figure 1. As indicated in Table 1, it contains the largest number of total Q&As and interaction-specific Q&As, supporting a wide range of applications including scene description, prediction, and planning. Additionally, to facilitate multi-modal learning across motion, language, and vision modalities, for each motion data, we render both the corresponding bird’s eye view and front view images by the MetaDrive simulator [24].
Based on WOMD-Reasoning, we build our Motion-LLaVA upon LLaVA [28], which can accurately answer driving-related questions. To comprehensively absorb detailed motion information, we further replace LLaVA’s vision encoder with a motion vector encoder [51], a motion prediction model. With this design and a set of effective training and inference strategies ablated, Motion-LLaVA achieves high accuracy in both driving scenario understanding and interaction analysis, outperforming baselines by large margins. We further benchmark several settings including multi-modal model fine-tuning strategies on WOMD-Reasoning to show the advantage of Motion-LLaVA. It unleashes the analytic abilities of motion language models to provide in-depth descriptions and reasoning on driving scenarios.
Our contributions can be summarized as follows:
  • We propose WOMD-Reasoning, the largest multi-modal dataset with 3 million Q&A pairs centered on interaction and intention reasoning in driving. It provides extensive insights into critical but previously overlooked interactions induced by traffic rules and human intentions.
  • We introduce Motion-LLaVA, a motion-language model designed for WOMD-Reasoning with robust capabilities, by benchmarking task performance across various configurations, including varying input modalities (vision v.s. motion), reasoning techniques (single step v.s. chain-of-thought), and network architectures (motion encoder pre-training).
  • Fine-tuned with WOMD-Reasoning, Motion-LLaVA is capable of providing detailed and insightful descriptions and interaction reasonings on various driving scenarios. Extensive evaluations with various metrics validate its wide coverage and high quality, especially in reasoning traffic rule-induced interactions.
Table 1: Comparison between WOMD-Reasoning Dataset and Previous Real-world Language Datasets for Driving Scenarios. Compared to previous datasets, WOMD-Reasoning covers comprehensive kinds of interactions, with a significantly higher size supporting a wide range of potential applications.
Dataset Data Source Statistics Interactions Applications
Total Scenes Total Q&As Interaction Q&As Distance-induced Traffic Rule-induced Human Intention-induced Scene Descriptions Motion Prediction Motion Planning
nuScenes-QA[37] nuScenes[5] 34k 460k 0
BDD-X[20] BDD[56] 26k 26k 26k
DriveLM[45] nuScenes[5] 5k 443k $\approx$199k
DRAMA[30] DRAMA[30] 18k 102k <18k
WOMD-Reasoning WOMD [13] 63k 2,940k 409k
2 Related Work
Application of Language Modal in Autonomous Driving. In recent years, language has been widely explored in autonomous driving, as incorporating language into implicit features could enhance the explain-ability [59, 40, 46, 6, 31]. For example, Kuo et al. [22] employ language features in the trajectory prediction model, ensuring both explainability and consistency between language and predictions. Similarly, Zhong et al. [64] use language input and LLMs to generate conditions in diffusion models, thereby controlling the generation of driving scenarios. Recently, general driving agents have been developed [57, 7, 45, 32], aiming to integrate all functions—perception, prediction, and planning—into a single language agent. However, we observe that their performance in interaction analysis is not entirely satisfying. This shortcoming is largely attributed to the lack of interaction analysis in language datasets specific to driving scenarios.
Language Datasets in Autonomous Driving. Language datasets for driving scenarios have recently been developed to support related LLM-based work. Due to the controllability of simulations, several simulation-based datasets have been created [7, 45]. However, simulations often fail to capture real-world interactions. To address this, recent studies [20, 41, 30] incorporate human labeling of real-world driving scenarios to ensure the inclusion of genuine interactions. Nevertheless, obtaining high-quality human labels is labor-intensive and costly, limiting the size and coverage of these datasets. Moreover, interactions driven by traffic rules and human intentions have often been overlooked. To streamline the labeling process, some studies [10, 33, 37, 45] have generated rule-based labels, although these typically cover only basic language elements like scene descriptions, leaving interaction analysis to human labelers. To minimize human labor, we first employ manual rules and then utilize ChatGPT-4 [34] to build our dataset automatically. The resulting dataset, WOMD-Reasoning, is rich in interaction details, particularly those stemming from traffic rules and human intentions. The designed automated data-curation pipeline enables the dataset to be significantly larger. We also create a high-quality human-verified subset. Preliminary human evaluation provides positive feedback with an approximate accuracy of 90%.
Fine-tuning LLM for Driving Tasks. Recent years have witnessed a surge of research on leveraging LLMs and vision-language models (VLMs) for driving applications [43, 7, 19, 58, 45, 44, 31, 47]. However, handling heterogeneous inputs such as vectorized motion representation remains challenging due to their domain gaps to most existing VLMs and LLMs. Notable prior work, such as [31], designed natural language prompt structures to accommodate numerical motion data, while another approach [7] proposed training LLMs with custom-built motion vector encoders.
Our Motion-LLaVA approach is inspired by the design of [7], augmenting existing LLM architectures with tailored components. In our study, we scale the training dataset by approximately 20 times and use real-world scenarios instead of simulation-based data facilitated by WOMD-Reasoning. Furthermore, we employ a motion prediction model as a motion vector encoder, utilizing its prior knowledge to streamline the training pipeline. To enhance reasoning accuracy, we adopt a Chain-of-Thought (CoT) [55] approach, allowing the model to base its reasoning on factual grounding, mitigating potential hallucinations. With WOMD-Reasoning, our Motion-LLaVA demonstrates the practicality of applying VLMs to real-world vectorized motion data to accurately answer diverse driving-related questions correctly.
3 Method
3.1 Building WOMD-Reasoning Dataset
Manually constructing autonomous driving datasets with fine-grained natural-language labels regarding the interactions among road participants would involve intense human labor, which is one of the main reasons why previous datasets have insufficient interaction analysis and a very limited overall size [42]. To reduce human labor while obtaining reasonably useful data on vehicle interactions, we propose a fully automatic data-curation pipeline to label the WOMD dataset with language: We first develop a rule-based program to interpret the motion dataset, which contains trajectories and the HD map, into language; and then build a set of prompts to utilize the reasoning abilities of ChatGPT-4 [34] to generate interaction analysis and reasoning, and organize the results into the target Q&A format. Details of each step are attached in the Appendix A.1. Codes for interpreting the motion dataset as well as the full prompts for ChatGPT-4 are available at https://github.com/yhli123/WOMD-Reasoning.
Our analysis process is performed on Microsoft Azure ChatGPT-4-Turbo API, which costs around 12,750 USD. The details of WOMD-Reasoning will be provided in Part 4.
3.2 Motion-LLaVA: an Multi-modal Fine-tuning Method on WOMD-Reasoning
To showcase the application of WOMD-Reasoning dataset, we perform a set of multi-modal fine-tuning experiments on it to build our Motion-LLaVA. LLaVA [28] is chosen as the pre-trained multi-modal model, in order to accommodate motion data input and text prompts input at the same time, where the vision encoder should be adapted to accommodate motion data inputs. Previous work [7] directly attended vectorized motion data with learnable queries to aggregate motion information. Inspired by [44], which trained its vision encoder with future trajectories as supervision, we utilize encoders from motion prediction models for encoding motion data. Specifically in our work, we adopted the encoder component of the motion prediction model MultiPath++ [51] as the motion vector encoder for the LLM. Ablations in Table 5 prove the helpfulness of knowledge prior in this encoder for fine-tuning. We also ablate some detailed fine-tuning strategies in Appendix A.7. Our Motion-LLaVA fine-tuning pipeline, summarized as the Motion-LLaVA method, is illustrated in Figure 3.
During training, we unfreeze all components including the motion vector encoder, and train on all fields of Q&A pairs in WOMD-Reasoning simultaneously to avoid the potential catastrophic forgetting. To prevent Motion-LLaVA from forgetting its inherent LLM’s knowledge, we mix up a few questions irrelevant to the current driving scenario into the training process. [61] During inference, we adopt Chain-of-Thought (CoT) Reasoning strategy [55] to reduce the hallucination. Specifically, we let Motion-LLaVA first answer the facts questions, which include questions in map environment, ego, and other agents’ motion status. Then, we aggregate these answers and feed them into Motion-LLaVA as the contexts for answering reasoning questions, which refers to questions in interactions and intentions analysis. Examples of these contexts can be found in the Appendix A.5. We ablate in Part 5.4 that such a CoT strategy can effectively improve the quality of answering reasoning questions.
4 WOMD-Reasoning Dataset
In this section, we provide detailed information about WOMD-Reasoning. We first present the organization and statistics of the dataset, then provide quality analysis based on case studies and preliminary human evaluations. Finally, we compare our dataset to previous ones to show the strengths of WOMD-Reasoning in diverse interactions, huge size, and wide potential applications.
4.1 Dataset Overview
Table 2 provides the statistics of Q&As in WOMD-Reasoning dataset, as well as in every Q&A class. Our Q&As can be classified into 5 categories: The first 3 categories are Factual contents: Map Environment, which includes the existence and category of intersections, existence and places of stop signs and crosswalks, and count of lanes, etc. Ego Agent’s Motion Status and Other (Surrounding) Agents’ Motion status, which describe each agent’s speed, acceleration, direction, related traffic light and sign, and relative position to the intersection center as well as to the ego agent. The other 2 categories are Reasoning contents: Interaction, which includes a summary of interactions between each surrounding agent and the ego agent and the reasoning behind such interactions. Intention, which sums up all interaction information to give the predicted intention of the ego agent, considering its responses to all the interactions. The factual description of the map environment and agents’ motion status in the dataset are good for training language scene describers, while the interaction and intention reasoning parts can be used for training language-based prediction and planning agents.
We claim that to the best of our knowledge, WOMD-Reasoning is the largest real-world language dataset for driving. As shown in Table 1, we provide roughly 3 million Q&A pairs for 63k scenes in WOMD, which are 6-113 times greater than those of other datasets we compare to. Furthermore, WOMD-Reasoning contains 346k unique traffic rule-induced and intention-induced interaction Q&As, as well as 63k comprehensive intention predictions of the ego agent containing responses to the interactions, which rarely exist in previous datasets. We will show these features through examples and statistics in the following section.
Table 2: Summary of dataset split and the quantitative distribution across different categories in the WOMD-Reasoning dataset.
Set Map Ego Agent Other Agents Interactions Intentions Total Scenes
Training 188k 268k 1,635k 287k 52k 2,430k 52k
Validation 41k 58k 341k 59k 11k 510k 11k
Total 229k 326k 1,976k 346k 63k 2,940k 63k
Per Scene Ave. 3.58 5.09 30.88 5.41 0.98 45.95 -
4.2 Dataset Quality
To testify to the quality of WOMD-Reasoning, we show examples from it containing rich traffic rule-induced and human intention-induced interactions.
WOMD-Reasoning contains traffic rule-induced interaction information. Firstly, in Figure 1 (a), we show a traffic scenario of an intersection controlled by the stop sign. In this case, the agent #0 is quite far away from the ego agent at the starting (current) moment, thus their interactions would rarely be considered in previous datasets, as they mainly cover distance-induced interactions. However, the two agents do have a significant interaction: Controlled by the stop signs, the agent #0 has to yield to the ego agent, even though the ego agent is still far from the intersection. This case vividly proves the remarkable importance of the traffic rule-induced interactions we cover.
Another example is a traffic light-controlled intersection shown in Figure 1 (b). Similar to the previous one, agent #0 is initially not close to the ego agent. However, thanks to our analysis of traffic rules, our language dataset effectively captures the interaction that the agent #1 would yield to the ego agent who is turning left and having right-of-way due to an arrow green light. This further proves that traffic rules are essential in determining the interactions.
Figure 2: A demonstration of Q&As in each part of WOMD-Reasoning dataset and an example for human intention-induced interaction. We show Q&As in all categories regarding the scenario while demonstrating language analysis of overtaking, a human intention-induced interaction in WOMD-Reasoning dataset.
Figure 3: The Motion-LLaVA pipeline, which fine-tunes a multi-modal model with WOMD-Reasoning. Motion data go through pre-trained motion vector encoders from Multipath++ [51] and a projector layer to serve together with the questions in WOMD-Reasoning as the inputs for the LLM. The answers in WOMD-Reasoning serve as the supervision.
Table 3: Preliminary Human Evaluation of the Dataset. Note that we ask evaluators to comment on whether main part of the answer is correct, where minor errors would not be considered as wrong.
Q&A Class Map Ego Agent Other Agents Interactions Intentions
Total Q&As 38 57 339 57 10
Ave. ‘Accuracy’ 80.70% 95.32% 93.51% 94.73% 90.00%
Std. Error 6.08% 2.03% 3.08% 3.51% 10.00%
WOMD-Reasoning contains human intention-induced interaction information. Beside traffic rules, our dataset also covers human intention-induced interactions. We provide an additional example to show this, as well as to demonstrate a whole picture of each section of the dataset. In Figure 2, WOMD-Reasoning provides fruitful information on the human intention-induced interaction - overtaking, based on pattern recognition abilities built in the GPT. We also show 1-2 Q&As for each category of the dataset for completeness.
To grab a sense of the accuracy of WOMD-Reasoning, we also execute a preliminary human evaluation. Consent was obtained from all the people interviewed. Note that here we ask evaluators to think a Q&A pair as correct if the correct main answer to the question is included in the answer, so the main answer is reasonably based on the human evaluators’ interpretation of the traffic data. This also does not exclude minor errors (like errors in detailed numbers or very far-away agents), since such errors are hardly distinguishable by the human evaluators. We invite 3 people to judge all Q&As from 11 consecutive scenes of the dataset, which contains 501 Q&As in total. The results are shown in Table 3. We conclude that considering our dataset is fully automatically generated, it achieves a satisfying accuracy, or at least it would significantly reduce the human labor necessary to build a more reliable dataset later. Further human verification of the dataset is still ongoing.
4.3 Vision Extension with Simulations
To support applications requiring visual information, we provide an API to extend WOMD-Reasoning scenario descriptions to the vision domain, compensating for the absence of raw camera data in the original WOMD dataset. We achieve this by utilizing an open-source traffic scenario simulation platform, namely ScenarioNet [25], and employing the MetaDrive [24] simulator to replay WOMD-Reasoning scenarios in both BEV layout and ego-view images through 3D rendering. This approach enables visual inputs to support the application of WOMD-Reasoning for future research in the vision and language domain. For video generation, we produce a 90-frame 10 Hz video aligned with Waymo’s trajectory data, including 10 frames of historical footage and 80 frames of future trajectory. Users may choose to use the historical portion for prediction tasks.
4.4 Comparison to Existing Datasets
Comparing WOMD-Reasoning with existing datasets, as shown in Table 1, we are the largest real-world dataset in terms of the total scenes, total Q&As, and the total interactions included. Also, WOMD-Reasoning supports Q&As in scene descriptions, prediction, and planning. Together with the simulated video of each covered scenario, it is suitable for training LLMs to perform nearly all major tasks in autonomous driving. Diving even deeper into the interaction reasoning in these datasets, we summarize that most interactions in previous datasets are caused by a very near distance. However, as we show in Sec. 4.2, many interactions happen when the two counterparts are initially far away, but they establish interactions by following the traffic rules or by human intentions. Our WOMD-Reasoning contains a wealth of this interaction information, as shown in Figure 4, the vocabulary statistics of the interaction and intention part of our dataset.
5 Experiments with Motion-LLaVA
To demonstrate the benefits of WOMD-Reasoning in applications, in this section, we show the performances of Motion-LLaVA, a multi-modal model fine-tuned on WOMD-Reasoning. We show that the model gains significant abilities from WOMD-Reasoning in answering both factual and reasoning questions related to driving, which is supported by both qualitative showcases of its answers to driving-related questions and quantitative evaluations of language metrics.
5.1 Model Setup
We take LLaVA-v1.5-7b [28] as the pre-trained VLM. Since MultiPath++ [51] did not release their codes nor checkpoints, we pre-train the implementations by [21]. Our multi-modal fine-tuning takes 2 GPU days ($\approx$ 1 day on 2xNVIDIA A6000 GPUs) to train 1 epoch on the entire training set of WOMD-Reasoning. For evaluation, we randomly select $\approx$ 1,000 QA pairs from the validation set of WOMD-Reasoning.
5.2 Qualitative Results
We present a set of Q&A examples answered by Motion-LLaVA covering each category of WOMD-Reasoning in one specific driving scenario, shown in Figure 5. We observe that Motion-LLaVA model is capable of providing accurate narratives and analysis on diverse driving-related contents thanks to the wide coverage of WOMD-Reasoning.
As WOMD-Reasoning is proposed to focus on the interaction analysis, we further present a few interaction analysis Q&As by Motion-LLaVA in Figure 6, in various interaction scenarios. These results further confirm the capacity of Motion-LLaVA for analyzing interactions.
5.3 Benchmarks of Multi-modal Reasoning Models
To comprehensively benchmark our proposed Motion-LLaVA model and fine-tuned LLaVA, we utilize a suite of standard language metrics, including BLEU-4 [35], METEOR [3], CIDEr [52], SPICE [1], and ROUGE-L [27], along with GPT Score [38] to assess the quality of the generated content. Details on the prompts used to calculate the GPT Score can be found in Appendix A.3. We present evaluation metrics of all generated answers, as well as those in the two categories, i.e. the Facts which include the road environment and agents’ motion status, and the Reasonings which cover analysis on interactions and intentions. Benchmark results are shown in Table 4.
Figure 4: Selected vocabulary statistics of WOMD-Reasoning. We stat vocabularies strongly related to traffic rule-induced and human intention-induced interactions in the interaction and intention part of WOMD-Reasoning, illustrating that our dataset contains abundant such interaction descriptions and reasoning.
Table 4: Benchmarks of fine-tuned multi-modal models with language metrics.
Questions Model ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) GPT Score ($\uparrow$)
All Shuffled Ground-truths 0.481 0.311 0.263 1.68 0.375 1.48
Non-Fine-tuned LLaVA [28] 0.512 0.211 0.275 1.36 0.455 2.31
Fine-tuned LLaVA [28] 0.779 0.581 0.439 5.51 0.735 6.88
Motion-LLaVA (Ours) 0.792 0.616 0.449 5.69 0.744 7.02
Facts Shuffled Ground-truths 0.528 0.349 0.273 2.26 0.401 2.06
Only Non-Fine-tuned LLaVA [28] 0.561 0.322 0.326 1.50 0.483 2.46
Fine-tuned LLaVA [28] 0.821 0.702 0.495 6.05 0.783 6.91
Motion-LLaVA (Ours) 0.840 0.736 0.516 6.35 0.794 7.09
Reasonings Shuffled Ground-truths 0.479 0.326 0.281 1.81 0.462 1.87
Only Non-Fine-tuned LLaVA [28] 0.262 0.097 0.201 0.06 0.283 0.91
Fine-tuned LLaVA [28] 0.596 0.452 0.357 2.26 0.557 6.14
Motion-LLaVA (Ours) 0.614 0.474 0.366 2.52 0.571 6.76
Figure 5: An example set of answers to diverse driving-related questions generated by Motion-LLaVA on a WOMD scenario. We see that Motion-LLaVA is capable of answering all topics of questions covered by WOMD-Reasoning accurately.
We first verify the effectiveness of these language metrics with shuffled ground-truths, which attach shuffled true answers to questions randomly. Results confirm that low scores are granted for this model, indicating that a similar structure with wrong answer content would not be able to cheat our metrics. Then we benchmark the off-the-shelf LLaVA [28], where BEV plots are used as the vision input for the driving scenario. This non-fine-tuned version earns a very low score in all metrics, confirming that LLaVA itself cannot answer driving-related questions correctly.
Then, we perform fine-tuning on WOMD-Reasoning. We compare two models with different modalities as motion input: LLaVA using the BEV plot of the driving scene as the vision input [28], and Motion-LLaVA which uses motion data as input. We observe that both fine-tuned multimodal models significantly outperform previous baselines across all metrics and question types, underscoring the value of fine-tuning on WOMD-Reasoning for addressing driving-related queries. Compared to the fine-tuned LLaVA, Motion-LLaVA achieves superior performance across all metrics and question types, highlighting the advantages of utilizing the motion data as input. We hypothesize that this performance gap is due to the information loss that occurs when converting motion data into a BEV plot with a scale bar. We also visualize and compare the answers of Motion-LLaVA and these baselines on the same driving scene in Appendix A.4.
Figure 6: An example set of answers to interaction-related questions generated by Motion-LLaVA on various WOMD scenarios. We see that Motion-LLaVA is capable to perform diverse interaction reasonings.
In addition to the language metrics, we provide other objective metrics to evaluate answers to quantitative questions regarding Facts. Of the 1,000 questions in the factual test set, 130 are quantitatively assessable (e.g., position, speed, etc.). For these, we report the root median squared error (Median Error), and accuracy within a 1.0-meter tolerance (Acc1.0), following similar metrics used in [19] as shown in Table 6. Results confirm that Motion-LLaVA is capable of providing accurate answers to factual questions.
Table 5: Benefits of using pre-trained motion vector encoder for fine-tuning multi-modal models on WOMD-Reasoning for factual Q&As.
Motion Vector Encoder Pre-training ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) Acc1.0 ($\uparrow$) Median Error($\downarrow$) GPT Score ($\uparrow$)
None 0.814 0.684 0.491 5.95 0.774 30.0% 2.00 6.79
Motion Caption 0.824 0.711 0.498 6.05 0.775 42.6% 2.24 6.97
AutoEncoder 0.837 0.727 0.516 6.35 0.803 47.3% 1.41 7.07
Motion Prediction (Ours) 0.840 0.736 0.516 6.35 0.794 56.2% 1.00 7.09
Table 6: Accuracy of answers to factual questions.
Model Acc1.0 ($\uparrow$) Median Error ($\downarrow$)
Fine-tuned LLaVA [28] 10.9% 5.0
Motion-LLaVA (Ours) 56.2% 1.0
5.4 Ablation on the Design of Motion-LLaVA
Chain-of-Thought Inference. To address the hallucinated response problem [15, 49, 45] in reasoning interactions, we use a Chain-of-Thought (CoT) approach to provide Motion-LLaVA with structured context for more accurate reasoning. Table 7 compares model performance with and without CoT on interaction and intention analysis. We see that with CoT, the model performs better in GPT Scores. We further show that if our motion data translation program A.1.1 is used during inference to provide contexts, i.e., with factual description context (an example can be found in Appendix A.6) and CoT, the performance can be significantly improved, proving the potential of the CoT inference strategy.
Influence of Pre-training for the Motion Vector Encoder. To establish a pre-training strategy for the motion vector encoder, we examine factual question answering across Motion-LLaVA with four pre-training configurations: no pre-training, autoencoder-style pre-training, motion-caption pre-training [7], and pre-training with a motion prediction task. All configurations shared the same encoder structure. In the motion-caption pre-training, we pre-train all but LLM weights in Motion-LLaVA, by penalizing errors in captioning the motion. For autoencoder-style pre-training and pre-training with a motion prediction task, we retain only the encoder after pre-training, to function as the motion vector encoder for Motion-LLaVA.
The performances of models with these pre-training settings are presented in Table 5. Generally, pre-training motion vector encoders can improve the response quality. Specifically, pre-training with motion prediction tasks outperforms the other two approaches. We attribute this to the close alignment between WOMD-Reasoning ’s question categories and the information needed in motion prediction tasks. As a result, pre-training with the motion prediction task achieves the highest GPT Score and overall accuracy, especially in numerical precision reflected by $Acc_{1.0}$.
6 Conclusion
We first build WOMD-Reasoning, a language annotation dataset upon WOMD, and then we propose Motion-LLaVA, an LLM fine-tuned on it, both focusing on analyzing and reasoning interactions. Specifically, our WOMD-Reasoning emphasizes interactions induced by traffic rules and human intentions, such as interactions governed by traffic lights and signs, and interactions like overtaking and following, which are both rarely covered in previous datasets. Besides, WOMD-Reasoning provides 3 million Q&A pairs, which makes it the largest real-world language dataset in autonomous driving to the best of our knowledge, covering a range of topics from scene descriptions, prediction, to planning. Boosted by WOMD-Reasoning, we benchmark a set of methods, to propose Motion-LLaVA, which fine-tunes a multi-modal model to accurately answer questions covering both driving scenario descriptions and interaction reasonings. Motion-LLaVA proves the effectiveness of WOMD-Reasoning in assisting real-world driving through various language-involved downstream tasks.
Table 7: Effects of CoT in answering reasoning-related questions.
Reasoning Inference Method ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) GPT Score ($\uparrow$)
Direct 0.616 0.474 0.365 2.54 0.578 6.59
CoT 0.614 0.474 0.366 2.52 0.571 6.76
CoT w/ Facts Contexts 0.664 0.542 0.403 3.10 0.631 7.64
Discussions. Although WOMD-Reasoning demonstrates its effectiveness in case studies, human evaluations, and multi-modal fine-tuning, we only manually verified a small subset. With human verification ongoing, the dataset will achieve even greater reliability. Also, WOMD-Reasoning is subject to the limitations of WOMD, e.g. its accuracy and lack of vision modal. We leverage a simulation approach to address the lack of support for video-Q&A, while acknowledging that it still exhibits a domain gap from real-world visual input. As WOMD continues to release additional modalities such as camera features and lidar data recently, and as the iterative development of the simulator [25, 24] we use, these limitations are expected to be naturally addressed.
7 Acknowledgements
This work was supported in part by Berkeley DeepDrive, NSF Grant. 2235013, and the Microsoft Accelerate Foundation Models Research Initiative.
Appendix A Appendix
In this appendix, we first show the building process of WOMD-Reasoning in detail in A.1. Specifically, we show the construction of the interpreted language of the motion data with an example (A.1.1), and we revisit the detailed methods utilizing GPT to build the language dataset as well as providing the full set of prompts (A.1.2). Some examples of the vision modal of WOMD-Reasoning provided by simulations are presented in A.2.
Then, for Motion-LLaVA, we show the prompt for obtaining the GPT Score used for evaluation (A.3). We further provide qualitative comparison of benchmarked multi-modal models listed in Figure 5 (A.4). We also provide an example of aggregated context (A.5) as well as factual description context (A.6) used in Chain-of-Thought inference ablations for reference. Finally, an ablation on details in multi-modal model fine-tuning strategy is attached (A.7).
A.1 Details in Building WOMD-Reasoning
In this appendix, we show the details of how we build WOMD-Reasoning dataset, including detailed explanations on each step we take, the example of the translated motion scenarios, and the GPT prompts. The full version of codes and prompts used herein can be found at https://github.com/yhli123/WOMD-Reasoning.
A.1.1 Automatic Translation of Driving Scenarios into Language Descriptions
ChatGPT cannot effectively process raw data from motion datasets, as it is mainly trained on textual information rather than motion data. Therefore, we first translate the motion information into textual descriptions by a rule-based program. Accurate motion translation is critical to the quality of the dataset, as it serves as the only input observable to GPT. Ensuring thorough and well-structured translations is essential for GPT to perform accurate interaction analysis and reasoning.
To ensure enough information for analyzing interactions involving the ego agent, we translate the motion data into an ego-centric description of the traffic scenario. Specifically, before describing the ego agent status, we first narrate the map environment around the ego agent, including information about related intersections, lanes, stop signs, crosswalks, speed bumps, etc. The above information provides a comprehensive spatial guidance that facilitates describing the motion of either ego or other agents accurately. Then, we describe the status of the ego agent in the scenario, including its motion status (e.g., the velocity, acceleration, etc.) and the positional status (e.g., the lane it occupies and its position related to the intersection center). Furthermore, the traffic light and signal information related to it are summarized in the description. Besides the ego agent, the depiction of other agents, including other vehicles, bicycles, and pedestrians, is also crucial for GPT to analyze the interactions. Therefore, we then describe every other agent in the scenario in the same way as we describe the ego agent. We additionally include the positional relation between the other agent and the ego one to support the identification of every agent’s position.
Utilizing the methods above, we can translate motion datasets into languages. Due to the limitations of computational resources of GPT, we only choose to translate those scenarios that have objects_of_interest labels in WOMD, which indicates confirmed significant interactions happening in the scenario. We build two subsets of WOMD-Reasoning with the same setting: the training set is built on the training set of WOMD, while the validation set is built on the interactive validation set of WOMD. In total, we translate 63k scenarios into language. An example of the translation can be found in Table 8.
Descriptions on map environment: The ego agent is heading towards intersection. The intersection center is 18.0 meters in front of the ego agent, and is 17.0 meters on the left of the ego agent. The intersection is a 4 way intersection. Descriptions on ego agent: The ego agent is on the 1 lane from the right, out of 3 lanes. Its current speed is 8 m/s. It is accelerating. Traffic Light for the ego agent is red. The ego agent is approaching a crosswalk 3 meters ahead. Descriptions on surrounding (other) agents: Surrounding agent # 0 is a vehicle. It is 23 meters in front of the ego agent, and is 17 meters on the left of the ego agent. It is heading right of the ego agent. Its current speed is 6 m/s. It is accelerating. It is in the intersection. It is 29 meters away from the ego agent. It is approaching a crosswalk 3 meters ahead. Surrounding agent # 2 is a vehicle. It is 11 meters on the right of the ego agent, and is 0 meters behind the ego agent. It is heading left of the ego agent. It is not moving. It is on the same side of the intersection as the ego agent. It is 34 meters away from the intersection center. $\cdots$ Descriptions on agents after 3 seconds: The following is the description of the ego and surrounding agents after 3.0 seconds: Surrounding agent # 0 will be 8 meters in front of the ego agent, and will be 5 meters on the left of the ego agent. It will be heading right of the ego agent. Its speed will be 7 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 18 meters away from the intersection center. The ego agent will be 13 meters in front of the current place, and will be 2 meters on the right of the current place. It will be heading in the same direction as the current moment. Its speed will be 3 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 20 meters away from the intersection center. Surrounding agent # 2 will be 13 meters on the right of the ego agent, and will be 9 meters behind the ego agent. It will be heading left of the ego agent. It will not be moving. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 34 meters away from the intersection center. $\cdots$
Table 8: An example of the language translation of motion data.
A.1.2 GPT-based Interaction Analysis
With the language translation of WOMD as input, we further design a set of prompts to utilize ChatGPT-4 [34] to build the Q&A dataset, taking advantage of its abilities in analyzing interactions. The prompts come with 4 main parts. (1) System Prompt, which lets the GPT know its role and the format of its input, i.e. the language translation. (2) Responsibility Prompt, which describes the responsibilities of the GPT, including the questions it should ask and answer, as well as the output format it should use. This will be talked about in detail in the next paragraph. (3) Rules Prompt, which contains global rules to guide the output and the analysis. (4) In-context Prompt, which provides a few human-written examples of input-output pairs to perform in-context learning. An example of each part of the prompts can be found in Table 9, 10, 11, and 12 respectively.
The questions in the responsibility part of the prompt decide what would be present in the generated language Q&A dataset. To provide a comprehensive language dataset, and to present interaction analysis as accurately as possible, we organize the questions into a chain to enable a well-designed Chain-of-Thought [55, 55, 8]: First, we ask questions about the map environment, the ego and other agents’ motion. These questions would let the GPT provide Q&As useful for fine-tuning scene description agents while enhancing the GPT’s understanding of the input information at the same time. Then, we ask questions about the interactions that occur in the observable future period. To include traffic rule-induced and intention-induced interactions, we ask the GPT to think about the interaction between each pair of agents by answering a sequence of questions: Q1: Are the two agents vehicles, and are they close to each other with traffic rules or signals governing their movements? If the answer is yes, we request that GPT analyze their yielding relations according to the traffic rules. Otherwise, we ask Q2: Are the two agents vehicle and pedestrian respectively with intention conflicts? If so, the vehicle generally should yield to the pedestrian for safety. If both questions do not help find the interaction relations, we come to Q3: Does the scene show patterns of certain intention-induced interaction? We provide descriptions of a few common interactions like overtaking, following, etc. GPT can choose the best fit from these patterns to try to cover more interaction types. Based on this pipeline, we can provide interaction analysis with rich information on traffic rule-induced and human intention-induced interactions. After that, we can finally wrap up the interactions between the ego and all other agents to provide the intention of the ego agent.
System prompt for GPT:
You are an AI assistant to analyze real-world driving scenes. You are provided with the detailed information (formated as: [start of the input] the detailed information [end of the input]) of the real-world driving scenario, which primarily includes information about the ego vehicle (denoted as ’Ego Agent’), other vehicles, cyclists or pedestrians (denoted as ’Surrounding agent #n’), intersections and road lanes. Note that the input includes the current situation as well as the situation in a future moment for your reference. Unless you are told to guess, only provide information you are certain according to the input. The input is complete. Any agents or other things not mentioned in the input does NOT exist, and you should NOT guess what would happen if they exist.
Table 9: System prompt for GPT.
Responsibility Prompt for GPT:
You are responsible for the following responsibilities: <First>, print "[Env QA]". Generate Q&A on all the scenario description inputs about the scenario environment. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" The Q&A should cover ALL information about the CURRENT moment in the input description. No information in the provided future moment should be used. Your questions should include all the questions in the following categories except the answer is not mentioned in the input: (1) About the intersection: its existence? Its type? (2) About the lanes: How many lanes on ego’s side? (3) About stop signs: its existence? where is it? which direction is it for?
$\cdots$
<Second>, print "[Ego QA]". Generate Q&A on all the scenario description inputs about the the ego agent. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" The Q&A should cover ALL information about the CURRENT moment in the input description. No information in the provided future moment should be used. Your questions should include all the questions in the following categories: (1) About its motion: its speed? its motion status? (2) About its position: its position? its lane? (3) About its direction: its facing direction? its turning direction?
$\cdots$
<Third>, print "[Sur QA]" and generate Q&A on all the scenario description inputs about surrounding agents. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" The Q&A should cover ALL information about the CURRENT moment in the input description. No information in the provided future moment should be used. For each surrounding agent, your questions should include all the questions in the following categories except the answer is not mentioned in the input: (1) About itself: its type? (2) About its motion: its speed? its motion status? (3) About its position: its position to the ego agent? its position to the intersection? its lane?
$\cdots$
<Fourth>, start a new paragraph and print "[Int QA]". For each surrounding agent, start a new paragraph and provide one Q&A for the interactions between it and the ego agent. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" Here [Question content] should have the meaning "What interactions will happen between surrounding agent #n and the ego agent?" without any additional information like "considering what" or "given what", but you must use diverse ways to put it. (i.e. ask in different ways for different surrounding agents) When providing [Answer content], you should use the descriptions for both the current moment and the future moment, and analyze based on them. Pay attention to the special situations where you should consider that there is an interaction:
$\cdots$
<Fifth>, for the ego agent, start a new paragraph, print "[Intention]", and provide one Q&A for the intention of the ego agent. The Q&A must be in the following format: "[Q][Question content] [A][Answer content]" Here [Question content] should have the meaning "What will be the intention of the ego agent?", but you must use diverse ways to put it. When providing [Answer content], and think about following questions in the time period from now to the provided future moment based on given description: (1) What is the ego agent’s intention if no other agents exist? (2) If applicable, what actions does the ego agent take to respond to traffic lights, stop signs, crosswalks, speed bumps and traffic rules? If some traffic controls are not applicable, do not output on that. (3) In each interaction involving the ego agent, what actions do the ego and surrounding agent take to respond to each other? If one agent does not have interaction with the ego agent, exclude it in your answer.
$\cdots$
Remember to use diverse ways to put the [Answer content]. Do not include any numbers regarding speeds or positions in your [Answer content].
Table 10: Responsibility prompt for GPT.
Global rules prompt for GPT: Here are some global rules that your need to follow: 1. Please provide the summary and analysis of the driving scenario according to your responsibilities. 2. Never use bullet points. Be concise and compact. 3. Give natural language, no parenthesis. 4. Give the results in the order of the responsibilities. Always start a new paragraph for each responsibility, but one responsibility can take several paragraphs. Do not combine multiple responsibilities together. Finish one responsibility before starting another. 5. Do not show the intersection number, lane number or the ego agent’s number. The surrounding agents’ numbers need to be shown. 6. Output in full. Never omit anything regardless of any reason. You are allowed to give very long outputs. 7. Remember that all the agents and traffic controls in the scene have been provided, so if any agents or traffic controls are not mentioned in the scene, they do not exist. You don’t need to consider them or give any output about them. Therefore, do not output words like "not mentioned" because this is inaccurate. 8. The traffic light info is for the current moment only. The future traffic light info is unknown. 9. Any action caused by traffic controls like lights, stop signs or crosswalks may also constitute yielding interactions. Traffic controls are only an indication of right-of-way, but any actions that are caused by the traffic controls are still considered interactions. For example, if a car is waiting for the red light or a stop sign while someone other cross the intersection with green light, the car is still considered yielding to the other car. 10. When analyzing the future, talk about the period between the current and future moment provided, not the two moments.
Table 11: Rule prompt for GPT.
Context prompt for GPT:
Here is one given example that you can refer to to help you better understand you responsibility <Fourth> and <Fifth>: For the input case of: [start of input] The ego agent is in intersection. The intersection center is 10.0 meters in front of the ego agent, and is 1.0 meters on the left of the ego agent. The intersection is a 4-way intersection. It is going straight. It is exiting the intersection. It ’s current speed is 6 m/s. It is accelerating. The ego agent is approaching a stop sign 1 meters ahead. There are 4 stop signs in the intersection. The ego agent is approaching a crosswalk 1 meters ahead. Surrounding agent # 0 is a vehicle. It is 4 meters on the left of the ego agent, and is 2 meters behind the ego agent. It is heading the opposite direction as the ego agent. Its current speed is 9 m/s. It is accelerating. It is departing from the intersection. It is on the same side of the intersection as the ego agent. It is 12 meters away from the intersection center.
$\cdots$ [end of input]
you may give the following analysis: [Int QA] [Q] What kind of interaction can be expected between the ego agent and surrounding agent #0 given their positions and directions? [A] Surrounding agent #0 will have no interaction with the ego agent as their intended paths have no conflicts.
$\cdots$ [End Int QA]
[Intention] [Q] What will the ego agent aim to do in the upcoming moments? [A] The ego agent intends to continue exiting the intersection. There is a stop sign on its side but it has already stopped and started afterwards. Surrounding agent #1 will yield to it so the ego agent does not need to respond. Surrounding agent #3 will follow it but the ego agent does not need to respond as well. Therefore, the ego agent will continue in the intersection and finish its left turn. [End Intention]
Here is another given example that you can refer to to help you better understand you responsibility <Fourth> and <Fifth>: [start of input] The ego agent is heading towards intersection. The intersection center is 28.0 meters in front of the ego agent, and is 3.0 meters on the left of the ego agent. The intersection is a 3 way intersection. The ego agent is on the 2 lane from the left, out of 3 lanes. It ’s current speed is 0 m/s. It is decelerating. Traffic Light for the ego agent is red 9.6 meters ahead. The ego agent is approaching a crosswalk 3 meters ahead. Surrounding agent # 0 is a vehicle. It is 29 meters in front of the ego agent, and is 4 meters on the right of the ego agent. It is heading left of the ego agent. Its current speed is 6 m/s. It is accelerating. It is in the intersection. It is 29 meters away from the ego agent. The ego agent is at a speed bump.
$\cdots$ [end of input]
you may give the following analysis: [Int QA] [Q] What kind of interaction can be expected between the ego agent and surrounding agent #0? [A] The ego agent will yield to surrounding agent #0 because they are both approaching the intersection and the traffic light is red for the ego agent, indicating that the ego agent must stop. [End Int QA]
[Intention] [Q] What will the ego agent aim to do in the upcoming moments? [A] The ego agent intends to enter the intersection, but it has to stop at the red traffic light and wait for it to change to green before proceeding through the intersection. It will also yield to surrounding agents #0 who is already in the intersection. Therefore, the ego agent will wait until the lights turn green and the surrounding agent #0 passes before entering the intersection. [End Intention]
$\cdots$
Table 12: In-context prompt for GPT.
Figure 7: More visual examples of WOMD-Reasoning through simulations.
A.2 Examples of Simulated Vision Modality
Figure 7 provides some whole examples of the simulated scenarios generated by MetaDrive [24] simulator using ScenarioNet [25] traffic scenario simulation platform.
A.3 Prompts for evaluating GPT Score
We employ the OpenAI GPT-4-turbo model [34] to evaluate the performance of all models benchmarked on the WOMD-Reasoning. The GPT Score prompt is presented in Table 13. Notably, we design our GPT Score prompt using a similar template to the one from [7].
We would like to request your feedback on the performance of an AI assistant in response to the user question displayed above. The user asks the question on observing a driving scenario. The correct answer is included for your reference. ### Scoring rules: Your scores for each answer should be between 0 (worst) and 10 (best). If the answer is mostly incorrect but includes important relevant information, such as identifying correct agents or actions, give a score of less than 3. Give intermediate scores for partially correct answers or when the numbers are close. Only give a score of 10 for flawless answers. Think critically and carefully. Most importantly, check the answer for factual correctness, but reward partial credit where applicable. ### Grading process For each of the xxx questions, provide a one line assessment of the student’s answer in the format: “‘ n. <short one sentence assessment>. Score: <score 0-10>. “‘ A few examples of what a good assessment might look like: 1. Acknowledged question unrelated to driving and attempted to answer. Score: 10. 2. Doesn’t answer the question but has all information necessary. Score: 3. 3. Incorrectly stated there are no vehicles around even though there is one. Score: 0. 4. Unable to answer question given information available. Score: 10. 5. Give 12 m/s or m or percentage for the correct answer of 13 (within 10% deviation). Score: 9. 6. Give 15 m/s or m or percentage for the correct answer of 13 (within 20% deviation). Score: 7. 7. Give 18 m/s or m or percentage for the correct answer of 13 (within 40% deviation). Score: 5. 8. Give 2 m/s or m or percentage for the correct answer of 13. Score: 0.
Table 13: Prompt for GPT Score.
Table 14: Refinement of Motion-LLaVA training strategy.
Multi-modal Fine-tuning Strategy Projector Alignment Encoder Freezing ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) Acc1.0 ($\uparrow$) Median ($\downarrow$) GPT Score ($\uparrow$)
LLaVA 0.821 0.678 0.484 5.82 0.771 2.3% 12.00 6.17
$\bm{-}$ Projector Alignment 0.821 0.697 0.492 5.96 0.782 2.3% 8.06 6.49
$\bm{-}$ Encoder Freezing (Ours) 0.840 0.736 0.516 6.35 0.794 56.2% 1.00 7.09
A.4 Qualitative Comparison of Benchmarked Multi-modal Models on Answering Questions in WOMD-Reasoning
In this appendix, we present examples of answers to questions in WOMD-Reasoning generated by the benchmarked multi-modal models listed in Table 4. The selected scenario corresponds to Figure 5, which already showcases answers from Motion-LLaVA. Additionally, we provide results for both the Non-Fine-tuned (i.e., off-the-shelf) and Fine-tuned LLaVA models, both using BEV plots as vision input.
For the Non-Fine-tuned LLaVA, we perform prompt engineering to make the answers coherent. However, the low GPT scores indicate that it remains challenging for LLaVA to directly comprehend the scenario:
  • Q: Can you describe the type of intersection present in the current driving scenario?
    A: The type of intersection present in the current driving scenario is a four-way intersection.
    GPT Score: 0. Incorrectly identifies the type of intersection.
  • Q: What is the ego agent’s current motion status?
    A: The ego agent is currently stopped at the intersection.
    GPT Score: 0. Incorrectly describes the motion status of the ego agent.
  • Q: Where is surrounding agent #5 in relation to the ego agent?
    A: Surrounding agent #5 is located to the right of the ego agent.
    GPT Score: 0. Incorrect location of surrounding agent #5 in relation to the ego agent.
  • Q: What kind of interaction will occur between the ego agent and surrounding agent #5?
    A: The ego agent will pass by the surrounding agent #5 at the intersection.
    GPT Score: 0. Incorrect prediction of the interaction between the ego agent and surrounding agent #5.
Quantitatively, the Non-Fine-tuned LLaVA achieves a remarkably low GPT Score of 2.31 across all questions in the validation set, demonstrating difficulty in understanding the driving scenarios based solely on the prior knowledge of the Vision-Language Model (VLM). For comparison, we evaluate the performance of the Fine-tuned LLaVA with vision input under the same scenario. The fine-tuned model achieves much more reasonable results, as illustrated below:
  • Q: Can you describe the type of intersection present in the current driving scenario?
    A: The intersection is a 3 way intersection.
    GPT Score: 10. The student’s answer matches the correct answer exactly.
  • Q: What is the ego agent’s current motion status?
    A: The ego agent is accelerating.
    GPT Score: 10. The student’s answer matches the correct answer exactly.
  • Q: Where is surrounding agent #5 in relation to the ego agent?
    A: It is 14 meters on the right and 10 meters in front of the ego agent.
    GPT Score: 0. The student’s answer incorrectly describes the position of surrounding agent #5.
  • Q: What kind of interaction will occur between the ego agent and surrounding agent #5?
    A: Surrounding agent #5 will follow the ego agent as they are both heading towards the intersection and are on the same side.
    GPT Score: 5. The student’s answer is partially correct but does not specify that agent #5 is behind the ego agent.
The results indicate that fine-tuning significantly improves LLaVA’s ability to interpret driving scenarios. This improvement can be attributed to the inclusion of our WOMD-Reasoning dataset. However, despite these enhancements, notable under-performance persists in questions related to Other Agent Status and Interaction & Intention, especially when compared to the results in Figure 5, which are generated using Motion-LLaVA. We attribute this limitation to potential information loss during the conversion of vectorized motion data into a BEV plot with a scalar bar.
A.5 Example Aggregated Context for Chain-of-Thought Reasoning
We present an example of an aggregated context, derived from the answers generated by Motion-LLaVA for factual questions, which is used by Motion-LLaVA to infer answers to reasoning-based questions. This context corresponds to the CoT configuration in Table 7:
The ego agent is turning left. Its current speed is 5 m/s. It is accelerating.
Surrounding agent #7 is a vehicle. It is 26 meters on the left of the ego agent and 8 meters in front. It is heading right of the ego agent. Its current speed is 3 m/s. It is decelerating.
The blue and orange sections contain answers to questions appearing in WOMD-Reasoning: blue addresses the Road Environment and Ego Agent Status, while orange covers Other Agent Status.
To generate this context, we prompt the Motion-LLaVA model with relevant questions and then concatenate the answers to create a complete context.
A.6 Example Factual Description Context for Chain-of-Thought Reasoning
We provide an example of a factual description context, generated using our motion data translation program as part of constructing WOMD-Reasoning, as described in Appendix A.1.1. This context is used by Motion-LLaVA to infer answers to reasoning-based questions and corresponds to the CoT w/ Facts Contexts configuration in Table 7. Notably, these contexts exclude future information compared to the translated motion data, thereby avoiding any information leakage:
The ego agent is in intersection. The intersection center is 6.0 meters in front of the ego agent, and is 0.0 meters on the left of the ego agent. The intersection is a 4 way intersection. It is turning left. It is exiting the intersection. Its current speed is 3 m/s. It is accelerating. Traffic Light for the ego agent is green. The ego agent is approaching a crosswalk 3 meters ahead.
Surrounding agent #7 is a vehicle. It is 26 meters on the left of the ego agent, and is 8 meters in front of the ego agent. It is heading right of the ego agent. Its current speed is 4 m/s. It is decelerating. It is heading towards the intersection. It is on the left of the intersection. Traffic Light for it is red 14.5 meters ahead. It is 26 meters away from the intersection center. It is approaching a crosswalk 4 meters ahead.
Similar to the aggregated context in Appendix A.5, the blue and orange sections contain answers to questions appearing in WOMD-Reasoning: blue addresses the Road Environment and Ego Agent Status, while orange covers Other Agent Status.
A.7 Ablation on Motion-LLaVA Training Strategy
The original LLaVA [28] training pipeline consists of two stages: a projector alignment stage (also known as pre-training) to align vision and language features, followed by a fine-tuning stage to enhance instruction-following capabilities. Throughout this process, the vision encoder remains frozen. Since our Motion-LLaVA builds on the architecture of the original LLaVA [28], it is intuitive to adopt a similar training approach for the WOMD-Reasoning task.
However, based on extensive experiments with fine-tuning strategies in Table 14, we find that incorporating projector alignment stage or keeping the motion vector encoder frozen throughout, which are practices employed in LLaVA [28], resulted in degraded performance on WOMD-Reasoning. Consequently, our Motion-LLaVA adopts a single-stage fine-tuning approach, where the motion vector encoder is unfrozen.